### v1

@@ 6.2 Evaluation Metrics
In the training phase, mean squared error (MSE) is used as the cost function for both traffic prediction and burst event detection networks. In the bursty event prediction problem, we usually deal with a kind of imbalanced data consisting of two classes which one of them is a rare event (bursty event samples). In this condition, using MSE metric to evaluate the performance of the burst detection network cannot clearly demonstrate the accuracy of the predictions [56]. For example, in a case that $94\%$ of the data is “00” and only $6\%$ is “$1$”, a network which always returns “00” will show an accuracy of $94\%$ in terms of MSE error, but in fact has a terrible performance. On the other hand, a network that detects some of the “$1$”s correctly and achieves an MSE value of $92\%$, may provide more information than the first network. Thus, besides the MSE metric, three other metrics are used in this paper to show the performance of the burst detection network. The first metric is called precision and defined as [56]:
@@ 6.3 Simulation Results of Traffic Prediction Network
The training loss of the traffic prediction network is illustrated in Fig. 6, indicating the network’s convergence to a final value in less than thirty iterations. Figures 7 and 8 present the output patterns for $T_{pred}=1[sec]$ and $T_{pred}=2[sec]$, respectively, with buffer sizes of $100$ and $200$ time slots for the rolling method. In Fig. 7, the predicted streams by the FLSP algorithm almost perfectly match the actual streams. The evaluation loss for this configuration is $3.96$ in terms of MSE, representing a substantial improvement of $94.15\%$ over the traditional rolling method, which yields a loss value of $67.73$. Meanwhile, the predictions in Fig. 8, as expected, demonstrate slightly lower accuracy due to increased prediction time ($T_{p}$), but still demonstrate an acceptable level of performance. The evaluation loss for this configuration is $11.76$ for FLSP algorithm, signifying a $70.19\%$ improvement compared to the traditional rolling method.
@@
Table 4: Performance comparison of rolling method and FLSP algorithm. Input sequence in all cases is concatenation of buffered data and fresh data collected from the mMTC network.
@@ 6.4 Simulation Results of Burst Detection Network
Performance metrics of the training phase for $T_{pred}=1\ [sec]$ are depicted in Fig. 9. It is observable from Fig. 9 (a) that the MSE error remains fairly constant after $10$ epochs, while the F1-score and other new metrics improve till epoch $15$ and then stabilize. 5 summarizes the performance evaluation of burst detection network on the test data sequences. In these simulations, the output of the FLSP and rolling algorithms is fed to the burst detection network as bunches of data. The F1-score metric best illustrates the superiority of the FLSP algorithm in providing high-quality information to the burst detection network, resulting in an F1-score value of $0.92$ out of $1$ for $T_{pred}=1\ [sec]$ and $0.86$ for $T_{pred}=2\ [sec]$.
@@
Fig. 10 shows the predicted bursty regions and actual bursty regions for $T_{pred}=1$ and $2$ seconds. As indicated by the arrows in Fig. 10 (b), some regions are detected as congested by error when $T_{pred}=2[sec]$, whereas congested regions in 10 (a) are detected almost perfectly when $T_{pred}=1[sec]$. Fig. 11 illustrates few examples of perfect and erroneous predictions by the proposed framework. Fig. 11 (a) shows a perfectly predicted bursty region that fits to the expected output. The proposed network predicts the bursty region in Fig. 11 (b) almost a second later than the expected time. Fig. 11 (c) depicts an example of false detection by the proposed network in which a non-congested region is detected as congested region. Late detection and false detection are happen less when $T_{pred}$ is small. As discussed earlier, there is always a trade-off between simulation time, accuracy, and the amount of time available to the network controller.
@@ 7 Conclusion and Future Work
Future work will focus on optimizing the parameters of our method and extracting the equations and metrics for the design problem. We will also investigate the effectiveness of the FLSP algorithm when applied to other types of recurrent neural networks, such as ConvLSTM, with the goal of source traffic prediction. Additionally, given the superior ability of our proposed algorithm to predict time-series data accurately, we plan to apply it to new problems requiring precise predictions, such as predicting the age of information (AOI) in time-critical scenarios. By further exploring these avenues, we aim to enhance the performance and applicability of our ML framework and contribute to the advancement of mMTC network capabilities and efficiency.

### v2

@@ 4.3 Evaluation Metrics
In the training phase, MSE is used as the cost function for both traffic prediction and burst event detection networks. In the bursty event prediction problem, we usually deal with a kind of imbalanced data consisting of two classes which one of them is a rare event (bursty event samples). In this condition, using MSE metric to evaluate the performance of the burst detection network cannot clearly demonstrate the accuracy of the predictions [51]. For example, in a case that $94\%$ of the data is “00” and only $6\%$ is “$1$”, a network which always returns “00” will show an accuracy of $94\%$ in terms of MSE error, but in fact has a terrible performance. On the other hand, a network that detects some of the “$1$”s correctly and achieves an MSE value of $92\%$, may provide more information than the first network. Thus, besides the MSE metric, three other metrics are used in this paper to show the performance of the burst detection network. The first metric is called precision and defined as [51]:
@@ 4.4 Simulation Results of Traffic Prediction Network
The training loss of the traffic prediction network is illustrated in Fig. 7, indicating the network’s convergence to a final value in less than thirty iterations. Figures 8 and 9 present the output patterns for $T_{pred}=1[sec]$ and $T_{pred}=2[sec]$, respectively, with buffer sizes of $100$ and $200$ time slots for the rolling method. In Fig. 8, the predicted streams by the FLSP algorithm almost perfectly match the actual streams. The evaluation loss for this configuration is $3.96$ in terms of MSE, representing a substantial improvement of $94.15\%$ over the traditional rolling method, which yields a loss value of $67.73$. Meanwhile, the predictions in Fig. 9, as expected, demonstrate slightly lower accuracy due to increased prediction time ($T_{p}$), but still demonstrate an acceptable level of performance. The evaluation loss for this configuration is $11.76$ for FLSP algorithm, signifying a $70.19\%$ improvement compared to the traditional rolling method.
@@
Table 6: Performance comparison of rolling method and FLSP algorithm when using LSTM for prediction. Input sequence in all cases is concatenation of buffered data and fresh data collected from the mMTC network.
@@ 4.5 Simulation Results of Burst Detection Network
Performance metrics of the training phase for $T_{pred}=1\ [sec]$ are depicted in Fig. 10. It is observable from Fig. 10 (a) that the MSE error remains fairly constant after $10$ epochs, while the F1-score and other new metrics improve till epoch $15$ and then stabilize. 7 summarizes the performance evaluation of burst detection network on the test data sequences. In these simulations, the output of the FLSP and rolling algorithms is fed to the burst detection network as bunches of data. The F1-score metric best illustrates the superiority of the FLSP algorithm in providing high-quality information to the burst detection network, resulting in an F1-score value of $0.92$ out of $1$ for $T_{pred}=1\ [sec]$ and $0.86$ for $T_{pred}=2\ [sec]$.
@@
Fig. 11 shows the predicted bursty regions and actual bursty regions for $T_{pred}=1$ and $2$ seconds. The burst detection network receives the output from the traffic prediction network every $T_{fresh}$ seconds, generating a binary output, either 00 or $1$, indicating whether congestion is expected within the next $l_{f}$ time slots. Consequently, for each set of $l_{f}$ time slots, the burst detection network provides a single value that represents the probability of congestion. For the purposes of visualization in Fig. 11 and 12, we replicate the burst detection network’s output $l_{f}$ times and plot the scaled results. The scaling is achieved by multiplying the output by $20$ to enhance clarity in the visual representation. As indicated by the arrows in Fig. 11 (b), some regions are detected as congested by error when $T_{pred}=2[sec]$, whereas congested regions in 11 (a) are detected almost perfectly when $T_{pred}=1[sec]$. Fig. 12 illustrates few examples of perfect and erroneous predictions by the proposed framework. Fig. 12 (a) shows a perfectly predicted bursty region that fits to the expected output. The proposed network predicts the bursty region in Fig. 12 (b) almost a second later than the expected time. Fig. 12 (c) depicts an example of false detection by the proposed network in which a non-congested region is detected as congested region. Late detection and false detection are happen less when $T_{pred}$ is small. As discussed earlier, there is always a trade-off between simulation time, accuracy, and the amount of time available to the network controller. For example, increasing the frequency of fresh data collection (decreasing $l_{f}$) can increase the resolution of the burst detection outputs and improves the accuracy of the detection at the cost of increased complexity.
@@
4.6 Comparative Analysis of LSTM, GRU, and CNN-1D Models
Table 8 compares the performance of the LSTM, GRU, and CNN-1D models in predicting mMTC network traffic, along with the number of parameters for each model. The simulation results indicate that the GRU model can achieve performance similar to that of the LSTM model, while requiring fewer parameters and offering faster execution times. For example, the GRU-DenseNet model contains $62,557,502$ parameters and takes $0:23:05$ to generate the output sequence for $100$ streams, whereas the LSTM-DenseNet model has $81,322,502$ parameters and takes $0:51:54$ to generate the same output sequence. Furthermore, applying the FLSP algorithm to the GRU model significantly enhances its performance compared to the rolling algorithm, demonstrating the effectiveness of the proposed FLSP algorithm in improving the GRU model’s performance.
In contrast, the CNN-1D model demonstrates poorer performance due to its inability to capture long-term dependencies. For instance, with a prediction time of $T_{pred}=1$ second and using the FLSP algorithm, LSTM and GRU models achieve MSE of $1.61$ and $2.01$, respectively, while the CNN-1D model records an MSE of $4.20$ with an input size is $200$ time slots. Moreover, when we use the prediction of LSTM and GRU models for detecting the congestion in up coming slots, we achieve a better performance than CNN-1D models. For example, the GRU model achieves an F1-score of $0.92$ with the FLSP algorithm, whereas the CNN-1D model only reaches an F1-score of $0.74$.
Although CNN-1D models are faster than LSTM and GRU models, their performance is inferior, and they lack flexibility. Unlike LSTM and GRU models, which can adapt to varying input sizes and prediction durations, CNN-1D models have fixed input and output sizes, necessitating a new model instance for each configuration. Consequently, LSTM and GRU networks are more favorable for network operators due to their adaptability and higher accuracy.
Table 8: Performance comparison of LSTM, GRU, and CNN-1D models.
Prediction Method Simulation Time Prediction Accuracy (MSE) Burst Detection Accuracy ($T_{p}=1[sec]$) Parameter Count
$T_{p}=1[sec]$ $T_{p}=2[sec]$ Precision Recall F1-score
LSTM (FLSP) $0:41:54$ $1.61$ $6.78$ $0.95$ $0.89$ $0.92$ $81,322,502$
GRU (FLSP) $0:23:05$ $1.61$ $6.30$ $0.95$ $0.91$ $0.93$ $62,557,502$
GRU (Rolling) ($l_{buff}=100$ [time slots]) $0:26:28$ $33.79$ $98.00$ $0.19$ $1$ $0.31$
GRU (Rolling) ($l_{buff}=200$ [time slots]) $0:30:14$ $8.33$ $23.58$ $0.28$ $1$ $0.43$
CNN-1D (Rolling) ($W=100$ [time slots]) $0:07:19$ $4.87$ $14.76$ $0.71$ $0.78$ $0.74$ $1,267,620$
CNN-1D (Rolling) ($W=200$ [time slots]) $0:07:18$ $4.20$ $12.11$ $0.75$ $0.82$ $0.78$ $1,657,620$
@@ 5 Conclusion and Future Work
Future work will focus on optimizing the parameters of our method and extracting the equations and metrics for the design problem. To better understand the inner workings of the proposed ML architecture and the effectiveness of each parameter and input feature on network performance, we will employ explainable deep learning techniques as discussed in [52]. We will also investigate the effectiveness of the FLSP algorithm when applied to other types of recurrent neural networks, such as ConvLSTM, with the goal of source traffic prediction. Additionally, given the superior ability of our proposed algorithm to predict time-series data accurately, we plan to apply it to new problems requiring precise predictions, such as predicting the age of information (AOI) in time-critical scenarios. By further exploring these avenues, we aim to enhance the performance and applicability of our ML framework and contribute to the advancement of mMTC network capabilities and efficiency.